Hi guys,
Can u guys help me with this problem.
The perimeter of a rectangular garden is 324ft. If it's length is twice it's width.
A. Write an equation for the peremeter of the garden.
B. Write the equation for the relation of the length and the width
C. What are the dimension of the garden.
D. Prove that the area of the garden is 5,832 square ft.

Answer by mananth(16946)

The perimeter of a rectangular garden is 324ft. If it's length is twice it's width.
A. Write an equation for the perimeter of the garden.
Perimeter = 2(l+w)
2(l+w)=324
but l=2w
2(2w+w)=324
2*3w=324
6w=324
B. Write the equation for the relation of the length and the width
l=2w
C. What are the dimension of the garden.
6w=324
/6
w=54
l=2w=108
D. Prove that the area of the garden is 5,832 square ft.
Area = l*w
A= 54*108
A= 5,832 sq.ft